Solve for the values of x for each of the following exponential equations and




inequalities.




1. 6𝑥+2 = 362𝑥−3




2. 5𝑥−1 = 625




3. 54𝑥 ≥ 253𝑥+2





1
Expert's answer
2022-02-24T13:06:19-0500

1.


"6^{x+2}=36^{2x-3}"

"x+2=2(2x-3)"

"3x=8"

"x=8\/3"


2.


"5^{x-1}=625"

"5^{x-1}=5^4"

"x-1=4"


"x=5"

3.


"5^{4x}\ge25^{3x+2}"

"4x\ge2(3x+2)"

"2x\le-4"

"x\le -2"